OpenStudy (kendricklamar2014):

Find the solution to the system of equations by using either graphing and substitution: Y=6-x and Y=x-2 A. (2,4) B. (-4, 2) C. (4,2) D. No Solutions Is the answer C?

OpenStudy (kendricklamar2014):

@beccaboo333 @Destinymasha @Eli_Moses @ganeshie8 @Hero @Reaper534

OpenStudy (anonymous):

are you in anon?

OpenStudy (anonymous):

no solution of these equations..!!! but can have if a tranformation eq from Y to y and X to x is given..!!!

OpenStudy (eli_moses):

6-x=x-2 -6 -6 -x =x+4 -x -x -2x=4 -------- -2 -2 x =-2 y=mx+b y=-2x+b can u do the rest from there or do you need more help??

OpenStudy (kendricklamar2014):

Im confused @Eli_Moses ?

OpenStudy (eli_moses):

which part

OpenStudy (kendricklamar2014):

y=mx+b and y=-2x+b

OpenStudy (eli_moses):

basically you're just substituting x and y in both equations to find the coordinate that can be the solution for both equations. C use this website to see it on a graph https://www.desmos.com/calculator just input both equations and find the point of intersection point of intersection=solution

OpenStudy (kendricklamar2014):

Ok thanks what's C^^

OpenStudy (eli_moses):

u'll find out

OpenStudy (kendricklamar2014):

@Squirrels

OpenStudy (kendricklamar2014):

@AccessDenied @agent0smith

OpenStudy (accessdenied):

You believe the answer is C? You can confirm whether it is or not by just plugging in the coordinate into each equation. C) (x, y) = (4, 2) y = 6 - x and y = x - 2 x=4, y=2 2 = 6 - 4 and 2 = 4 - 2. Are both of these equations true?

OpenStudy (kendricklamar2014):

No

OpenStudy (accessdenied):

Which one looks wrong? 2 = 6 - 4 simplify the right side 2 = 2 That is a true statement; we can't argue that 2 equals itself!

OpenStudy (kendricklamar2014):

yeah ur right

OpenStudy (accessdenied):

And for 2 = 4 - 2 We simplify 4 - 2 and get 2 as well. 2 = 2 again, and so (C) does work to solve this system!

OpenStudy (kendricklamar2014):

thanks

OpenStudy (accessdenied):

You're welcome!